[Hallc_running] [Revised Logentry] RC Update - 01/07/22

sjdkay at jlab.org sjdkay at jlab.org
Fri Jul 1 10:10:01 EDT 2022

Logentry Text:
Some pretty good news overall today. The work yesterday had us down until late swing *but* things are actually looking up following that.

- Following the work yesterday, MCC got us up to 75uA for more than just 2 seconds. See - https://logbooks.jlab.org/entry/4009273 https://logbooks.jlab.org/entry/4009266 https://logbooks.jlab.org/entry/4009261 https://logbooks.jlab.org/entry/4009254
- 70+uA was a bit too unstable, so we ran with 60uA instead.
- We got stable(ish) 60uA beam late swing, it was a bit too trippy so we dropped to 50uA overnight
- We lost an hour due to an SHMS dipole interlock at around 2am - https://logbooks.jlab.org/entry/4009339 - Steve L beamed himself in to fix it
- Following that we took 50uA beam overnight, as of 09:30 this morning, we had 38.5k/40k for our LD- setting - Q2=3.85, W=3.07 (left)
- At that point, we lost beam and it was going to be an hour or so till it's back. We're in the process of moving back to positive polarity running now, magnets are cycling and we're going to the Q2=6.00 W=2.40 x=0.55 θ_SHMS=16.92 (Left) setting on LD2
- MCC has some extra things they want to test too so it might be into the early afternoon before we get production beam back, the goal is stable 50-60uA beam through the long weekend

Other good news - No further HMS Q1 trips so far.

We'll have a meeting at 11:00, upstairs in the counting house and on zoom at - https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcEE0QT09


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4009479
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220701/fe1e2c07/attachment.html>

More information about the Hallc_running mailing list